# Multi-refined Feature Enhanced Sentiment Analysis Using Contextual Instruction

### v1

@@ 2.3 Instructing and Prompting
Figure 1: The framework of our proposed Model CISEA-MRFE.
@@ 3.1 Problem Definition
We define sentiment classification as predicting a polarity label $y\in\mathcal{S}=\{\text{positive},\ \text{negative},\ \text{neutral}\}$ for each review $r=(x_{1},\dots,x_{m})$, where each sentence $x_{i}=(w_{i,1},\dots,w_{i,n_{i}})$, and $w_{i,j}\in\mathcal{V}$. The review is first enriched with contextual instruction: $A=\mathrm{CI}(r)\in\mathbb{R}^{n\times d_{w}}$ where $n$ is the token length after augmentation and $d_{w}$ is the embedding dimension. Our model maps this to sentiment probabilities: $\hat{\mathbf{y}}=f_{\theta}(A)=\mathrm{CISEA\text{-}MRFE}(A)\in\Delta^{|\mathcal{S}|},\quad\hat{y}=\arg\max_{s}\hat{y}_{s}$ where $\Delta^{|\mathcal{S}|}$ is the probability simplex.
@@ 3.2 Context Instruction
where $\Delta^{3}$ is the 3-dimensional probability simplex corresponding to sentiment classes $\{\text{positive},\,\text{negative},\,\text{neutral}\}$.
@@ 3.2.1 Instruction Template Example
Comment: The review is positive because "fantastic" expresses strong sentiment.
@@ 4.1.1 Datasets
We utilize four real-world benchmark datasets to evaluate the performance of our CISEA-MRFE model. Table LABEL:tab:my_label_1 summarizes their statistics, including label distributions and average sentence lengths. All datasets are widely used in the sentiment analysis literature [14, 52, 30, 53].
@@ 4.1.3 Parameter Settings
Furthermore, we assess the impact of various kernel size combinations in the depth-wise CNN component, as detailed in Table 2. A sensitivity analysis is conducted to examine the configuration effects across datasets (see Section 5.2.6), with accuracy trends visualized in Figure 5. Full hyperparameter configurations are provided in Table LABEL:tab:8.
@@
Table 1: Dataset statistics
Dataset name Total Avg. Word/Sent. length class
@@ 4.2 Full hyperparameter configurations
Model configuration and training setup for CISEA-MRFE are shown in Table LABEL:tab:8.
Table 3: Model Configuration and Training Setup for CISEA-MRFE
@@
Max Kernel Size ($k_{\text{max}}$) 7
@@
Augmented Text Max Length 512
@@ 4.2.1 Baselines
Table 4: Comparison Results for different model types across the four datasets.
@@
Lexicon-based SAAT - - - - - - 87.96 87.94
SAWE - - - - - - 88.01 87.98
SAATWE - - - - - - 88.06 88.04
@@
Table 5: The results of various combinations of our methods across five PLMs, evaluated using Accuracy and Macro-F1 (M-F1) metrics, reveal no statistically significant differences between the two metrics based on t-statistics and corresponding p-values.
Model PLM Model IMDb Yelp Twitter Amazon
Accuracy M-F1 Accuracy M-F1 Accuracy M-F1 Accuracy M-F1
@@
tst.t-st. = 0.4443 tst.t-st.= 0.3390 tst.t-st.= -0.1050 tst.t-st. = 0.9623
pval.p-val.= 0.6686 pval.p-val.= 0.7433 pval.p-val.= 0.9190 pval.p-val. = 0.3641
@@
tst.t-st. = 1.9645 tst.t-st. = 0.6761 tst.t-st.= 0.8085 tst.t-st.= 0.6408
pval.p-val.= 0.0851 pval.p-val. = 0.5180 pval.p-val.= 0.4422 pval.p-val.= 0.5396
@@
tst.t-st. = 0.1042 tst.t-st. = 0.0759 tst.t-st.= 0.2829 tst.t-st.= 0.3049
pval.p-val. = 0.9196 pval.p-val. = 0.9419 pval.p-val.= 0.7844 pval.p-val.= 0.7682
Table 6: Ablation study on essential components of CISEA-MRFE across four datasets
@@ 4.3.4 Parametric Sensitivity (RQ5)
Figure 5: Comparative performance analysis of various multi-kernel configurations in the SADE module across four benchmark datasets. The bar plot illustrates the classification accuracy achieved by each kernel combination, while the overlaid line plot indicates the corresponding effective window size (maximum kernel span). The results suggest that integrating broader and more diverse receptive fields enhances local feature extraction, contributing to improved model performance across tasks
Figure 6: Performance of CISEA models on the four datasets.
@@ 4.4 Discussion
This study introduces CISEA‑MRFE, a comprehensive sentiment classification framework that integrates CI, SEA, and an MRFE module composed of SADE and EECE. By combining contextual guidance, paraphrastic augmentation, and multi‑granular feature refinement, CISEA‑MRFE is designed to address key challenges in sentiment analysis, including domain transferability, the modeling of nuanced emotional cues, and performance degradation under data imbalance. Empirical evaluations across four diverse benchmark datasets, including IMDb, Yelp, Twitter, and Amazon, demonstrate that CISEA‑MRFE consistently outperforms competitive baselines, underscoring its ability to effectively capture both short‑ and long‑form sentiment cues. Importantly, the model maintains high performance across multiple pre‑trained language model backbones (BERT, RoBERTa, BART, DistilBERT, GPT‑2), confirming its robustness and architectural transferability. These findings highlight the value of the CI module in disambiguating sentiment polarity under noisy settings, while SEA contributes to mitigating bias toward dominant sentiment classes through controlled semantic augmentation. Ablation analyses further reveal the critical role of the MRFE module, where SADE improves localized n‑gram sensitivity and EECE refines global emotional context. Removing these components results in significant performance degradation, confirming that multi‑scale feature refinement and emotion‑aware modeling are essential to the framework’s efficacy. Notably, even simplified CISEA‑MRFE variants surpass recent prompting‑based strategies, illustrating the superiority of explicit multi‑level feature integration over template‑only approaches. In summary, CISEA‑MRFE unifies instruction‑guided learning, emotion‑aware encoding, and multi‑scale feature extraction, delivering a scalable and versatile solution for real‑world sentiment analysis.

### v2

@@ 2.3 Instructing and Prompting
Figure 1: Overall framework of the proposed CISEA-MRFE model.
@@ 3.1 Problem Definition
We define sentiment classification as predicting a polarity label $y\in\mathcal{S}=\{\text{positive},\ \text{negative},\newline \ \text{neutral}\}$ for each review $r=(x_{1},\dots,x_{m})$, where each sentence $x_{i}=(w_{i,1},\dots,w_{i,n_{i}})$, and $w_{i,j}\in\mathcal{V}$. The review is first enriched with contextual instruction: $A=\mathrm{CI}(r)\in\mathbb{R}^{n\times d_{w}}$ where $n$ is the token length after augmentation and $d_{w}$ is the embedding dimension. Our model maps this to sentiment probabilities: $\hat{\mathbf{y}}=f_{\theta}(A)=\mathrm{CISEA\text{-}MRFE}(A)\in\Delta^{|\mathcal{S}|},\quad\hat{y}=\arg\max_{s}\hat{y}_{s}$ where $\Delta^{|\mathcal{S}|}$ is the probability simplex.
@@ 3.2 Context Instruction
where $\Delta^{3}$ is the 3-dimensional probability simplex corresponding to sentiment classes $\{\text{positive},\newline ,\text{negative},\text{neutral}\}$.
@@ 3.2.1 Instruction Template Example
Comment: The review is positive because "fantastic" expresses
strong sentiment.
@@ 4.1.1 Datasets
We utilize four real-world benchmark datasets to evaluate the performance of our CISEA-MRFE model. Table 1 summarizes their statistics, including label distributions and average sentence lengths. All datasets are widely used in the sentiment analysis literature [14, 52, 30, 53].
@@ 4.1.3 Parameter Settings
Furthermore, we assess the impact of various kernel size combinations in the depth-wise CNN component, as detailed in Table 2. A sensitivity analysis is conducted to examine the configuration effects across datasets (see Section 5.2.6), with accuracy trends visualized in Figure 5. Full hyperparameter configurations are provided in Table 3.
@@
Table 1: Statistics of the benchmark datasets used in this study.
Dataset Total Samples Avg. Words/Sentence Classes
@@ 4.2 Full hyperparameter configurations
Model configuration and training setup for CISEA-MRFE are shown in Table 3.
Table 3: Model configuration and training setup for CISEA-MRFE.
@@
Maximum Kernel Size ($k_{\text{max}}$) 7
@@
Augmented Text Maximum Length 512
@@ 4.2.1 Baselines
Table 4: Comparison results for different model types across four datasets.
@@
Lexicon-based SAAT - - - - - - 88.0 87.9
SAWE - - - - - - 88.0 88.0
SAATWE - - - - - - 88.1 88.0
@@
Table 5: Results of our method combinations across five PLMs on four datasets. Accuracy and Macro-F1 (M-F1) show no statistically significant differences based on $t$-statistics and corresponding $p$-values.
Model PLM IMDb Yelp Twitter Amazon
Acc M-F1 Acc M-F1 Acc M-F1 Acc M-F1
@@
$t$-stat. 0.4443 $t$-stat. 0.3390 $t$-stat. -0.1050 $t$-stat. 0.9623
pp-val. 0.6686 pp-val. 0.7433 pp-val. 0.9190 pp-val. 0.3641
@@
$t$-stat. 1.9645 $t$-stat. 0.6761 $t$-stat. 0.8085 $t$-stat. 0.6408
pp-val. 0.0851 pp-val. 0.5180 pp-val. 0.4422 pp-val. 0.5396
@@
$t$-stat. 0.1042 $t$-stat. 0.0759 $t$-stat. 0.2829 $t$-stat. 0.3049
pp-val. 0.9196 pp-val. 0.9419 pp-val. 0.7844 pp-val. 0.7682
Table 6: Ablation study on key components of CISEA-MRFE across four datasets.
@@ 4.3.4 Parametric Sensitivity (RQ5)
Figure 5: Comparative performance analysis of various multi-kernel configurations in the SADE module across four benchmark datasets. The bar plot illustrates classification accuracy achieved by each kernel combination, while the overlaid line plot indicates the effective window size (maximum kernel span). The results indicate that integrating broader and more diverse receptive fields enhances local feature extraction, improving model performance across tasks.
Figure 6: Performance of CISEA models across four benchmark datasets.
@@ 4.4 Discussion
This study introduces CISEA‑MRFE, a comprehensive sentiment classification framework that integrates CI, SEA, and an MRFE module composed of SADE and EECE. By combining contextual guidance, paraphrastic augmentation, and multi‑granular feature refinement, CISEA‑
MRFE is designed to address key challenges in sentiment analysis, including domain transferability, the modeling of nuanced emotional cues, and performance degradation under data imbalance. Empirical evaluations across four diverse benchmark datasets, including IMDb, Yelp, Twitter, and Amazon, demonstrate that CISEA‑MRFE consistently outperforms competitive baselines, underscoring its ability to effectively capture both short‑ and long‑form sentiment cues. Importantly, the model maintains high performance across multiple pre‑trained language model backbones (BERT, RoBERTa, BART, DistilBERT, GPT‑2), confirming its robustness and architectural transferability. These findings highlight the value of the CI module in disambiguating sentiment polarity under noisy settings, while SEA contributes to mitigating bias toward dominant sentiment classes through controlled semantic augmentation. Ablation analyses further reveal the critical role of the MRFE module, where SADE improves localized n‑gram sensitivity and EECE refines global emotional context. Removing these components results in significant performance degradation, confirming that multi‑scale feature refinement and emotion‑aware modeling are essential to the framework’s efficacy. Notably, even simplified CISEA‑MRFE variants surpass recent prompting‑based strategies, illustrating the superiority of explicit multi‑level feature integration over template‑only approaches. In summary, CISEA‑MRFE unifies instruction‑guided learning, emotion‑aware encoding, and multi‑scale feature extraction, delivering a scalable and versatile solution for real‑world sentiment analysis.